A square is inscribed inside a quadrant of a circle of radius 10 cm. Calculate the area of the square.
Answer by dabanfield(803)

The diagonal of the square is the radius of the circle. The diagonal of the square is also the hypotenuse of an isosceles triangle. Let x be the length of the sides. Then by the Pythagorean Theorem we have:
x^2 + x^2 = 10^2
2*x^2 = 100
x^2 = 50
But x^2 is also the area of the square.